how to do scalar product of vectors

Grade:6

4 Answers

naniuppati
10 years ago

abcosx

subha
10 years ago

vector A=3i+4j+5j, vector B = 1i+2j+3k           The scalor product of two vectors is = (3i+4j+5k)(1i+2j+3k)=3+8+15=26 this is the way to find scalor product of two vectors.

Janibasha
10 years ago

scalar product is also known as dot product. the formula is ABCOS$


if the two vectors be A andB then the scalarproduct of the two vectors is PRODUCT OF THE MAGNITUDE OF THE VECTORS and COSINE OF THE ANGLE BETWEEN THEM.










%%%%

G vamsinath reddy
10 years ago

ai+bj+ck*mi+nj+ok=ami+bnj+cok

here we product the i to i ,and j to j , c to c.